# BioCLIP 2: Emergent Properties from Scaling Hierarchical Contrastive Learning

### v1

@@ Appendix G Discussion on the Relationship with Neural Collapse
First, the class prototypes in standard classification tasks and supervised contrastive learning (SCL) can both be treated as fully trainable parameters [80]. In standard classification, the prototypes are weights of the linear classifier. In SCL, they are simply the mean embedding of the corresponding classes. The class prototypes will form a simplex ETF after extensive training. Comparatively, the adopted contrastive training scheme in BioCLIP 2 employs text embeddings of the taxonomic labels as class prototypes. During the creation of taxonomic labels, hierarchical structures have been naturally embedded into them based on ecological and functional evidence. Different species can share higher taxonomic levels, which makes them hard negatives. Therefore, even if the text encoder is also being trained, the generated prototypes will not become ETF.
@@ Appendix H Biological Visual Evaluation Details
Instead of training specialist models, we treat the evaluated models as frozen visual embedding extractors. Standard machine learning algorithms are applied on top of the acquired visual embeddings to predict the corresponding labels. Such a design is adopted to evaluate the quality of the embeddings while avoiding the influence of complicated optimization loops. In the following, we introduce the details of adopted benchmarks and the evaluation algorithms. All the experiments are conducted with 1 NVIDIA A100 GPU, and the running time for each task is within 30 minutes.
FishNet. FishNet focuses on recognizing, locating, and predicting species and their functional traits [16]. Specifically, images are collected with annotations of habitat, ecological role, and nutritional value. In this work, we mainly focus on the prediction of habitats and ecological roles, involving 9 groups of binary labels (e.g. whether the fish can live in freshwater). Following the practice in the original paper, we train a two-layer linear classifier with binary cross-entropy loss to predict the 9 labels. We count a correct prediction only if all the 9 labels are predicted correctly for the sample. The original train-test split is adopted, where images are used in training, and the remaining images are used for testing. This task evaluates whether the embedding distribution of different species is aligned with their ecological relationships.
@@ I.1 Taxonomic Standardization
When GNVerifier returns a result from a query, our taxonomic alignment package combines it with the input taxonomic hierarchy along with the query parameters to form a resolution attempt. The query response, based on the most specific taxonomic term available in the input data, along with the remaining input hierarchy (entry’s resolution attempt) is then matched against pre-defined profiles. The algorithm uses these three components to fit against the series of profiles to determine whether a confident resolution is found or if an alternative query strategy is needed (such as using a different query term or data source), iterating until a match is made or alternative approaches are exhausted. If all attempts ultimately fail to match a profile, the original input data is used for the entry’s label in the final output. After hierarchical resolution, common names are annotated. For each entry, the common name corresponding to its most specific resolved taxonomic term is selected, when available, using the preferred English vernacular names from the GBIF Backbone Taxonomy [85]. In the output, the proportion of records changed for each data source is EOL: %, FathomNet: %, BIOSCAN: %, and GBIF: %. The notably low modification rate for GBIF reflects our taxonomic alignment package’s preference for the GBIF Backbone Taxonomy as its primary reference. After standardization, we summarize the image number and taxa number from each source in Figure 12. We provide the source code for this part in TaxonoPy.
I.2 Image-quality Screens
At download, all images were checked to be sufficiently large ( pixels or more on the shortest side), and resized, where needed, so that the largest side does not exceed pixels. The code, support set class embeddings, and further details of the image-quality screens and the subsequent duplicate control are provided in our dataset repository, TreeOfLife-toolbox.
| Publisher (GBIF Data Source) | Distinct Taxa | Total Taxa | Images | % GBIF Taxa |
| iNaturalist.org | 81,671 | 350,229 | 134,877,019 | 9.07% |
| Natural History Museum | 38,535 | 148,637 | 3,603,256 | 4.28% |
| Museum national d’Histoire nat. | 37,201 | 258,766 | 6,185,477 | 4.13% |
| Naturalis Biodiversity Center | 25,421 | 244,697 | 5,282,428 | 2.82% |
I.2.1 Museum Specimen Processing
Museums often consider multiple specimens of the same species to be connected instead of differentiating them (as they are often collected from the same or similar location). Thus, these occurrences often include images of their metadata, duplicates of the same or different specimens under one occurrence ID, or less informative images. Some common complications to consider when working with museum specimen images that influenced our processing are that
1.
Plant and fungi specimens may be stored in envelopes or folders. These are often photographed and digitized under the same occurrence ID as the image of the specimen itself, thus creating extraneous images we do not want to include in training. They also sometimes pair this with living specimen images (which would be worth keeping as well). See Figure 13, which demonstrates variety within a single occurrence for fungi specimens. Plant specimens have a secondary confounding factor in that they are often pressed for preservation with their metadata on the page (see Figure 14).
2.
Fish, worms, and similar will be stored in jars. A single jar is considered an occurrence, so only one specimen may be photographed—perhaps at multiple angles—the jar may be photographed, multiple specimens photographed, etc. In any of these cases, the images will all be labeled with the same metadata that does not include this context.
3.
Fossilized specimens are often digitized with a photo of the fossil (appearing as a “rock”, lacking a clear view of the specimen) and a close-up of the specimen itself, e.g., in the case of a bug in amber (see Figure 15).
4.
For animal specimens, both of the above cases may occur: there may be simply a close-up of the tag (similar issue to envelopes with plants and fungi). There may also be multiple specimens photographed within the same occurrence (as noted with 2 about fish and worms). Museums often consider multiple specimens of the same species to be connected instead of differentiating them. This is two-fold, in that they are generally, collected from the same or similar location at or about the same time, and a specimen is kept as a representative of its species, so there is not a clear need for distinguishing between them. We also see multiple views of the same specimen within an occurrence. Examples in Figure 16.
In order to appropriately separate these images, we treated them as museums do, specifically by first dividing them into 11 collection areas (Fungi, Insect, Invertebrate Zoology, Microbiology, Plant, Uncategorized, and five classes of Vertebrate Zoology: Amphibians, Birds, Fishes, Mammals, Others) inspired by the Smithsonian Institution’s categorical subdivisions for their biological museum collections. From here, we further divided each category based on its image type (e.g., fossil or preserved specimen, as specified in GBIF metadata).
For each museum specimen category, we manually curated a small “support set” of representative specimen and non-specimen images. We embedded these examples using CLIP (ViT-L/14@336px) [7]; we chose not to use BioCLIP since museum specimen labels were not filtered from its training data (e.g., EOL contains many museum specimen images). To capture intra-class diversity, we ran K-Means on each support set and retained the resulting cluster centers. During classification, we L2-normalized both the input image’s embedding and each center, then assigned the image to the nearest center in Euclidean space. This processing was applied to all museum specimen images identified within GBIF. The support set embeddings are included in TreeOfLife-toolbox.
I.2.2 Camera Trap Images
Some occurrences include large-volume camera-trap sequences, with up to images. These occurrences have a single taxonomic label applied across all images, though there are different taxa (e.g., the first few frames may have a duck, while later images have a swan, another a goose, but the label for all is a duck). To reduce the risk of introducing such noise while still capturing relevant biodiversity, we filter the dataset to include only occurrences with 15 images or fewer. We then use MegaDetector [45, 46] to filter “empty” frames. In creating a dataset to train a foundation model on the entire tree of life, it is prudent to avoid introducing too many plant images labeled as animals.
I.2.3 Citizen Science Images
Similarly, citizen science occurrences may contain many images assigned a single label. For each occurrence, we thus embedded the images using BioCLIP [13], standardized them, calculated the pair-wise cosine similarity, and used the mean to indicate the occurrence’s “distinctness”. In these instances, we identified three broad categories into which to further subdivide them for reduction:
1.
Mixed occurrences: multiple species all labeled as just one of them (low similarity). These were treated as noise and discarded.
2.
Multiple image occurrences of same species: observations suggest these are creature close-ups and environment images under a single occurrence. They may also be larger groups with close-ups. For these, we randomly sample down to 5 images.
3.
Camera trap images: likely images collected in backyards (high similarity). Occasionally people upload images from camera traps to citizen science platforms like iNaturualist (also ex: MammalWeb, a citizen science platform for camera trap images). These are processed under camera trap protocols described above.
The final step in our quality control processing pipeline was to filter out identifiable images of people from the training data. This was done by running MTCNN [47] on all images downloaded from GBIF and EOL.
I.3 Duplicate and Leakage Control
GBIF and EOL are large biodiversity data aggregators, sourcing images that are also used in biological benchmarks such as iNat21 [35] and NeWT [18]. Due to metadata provenance complications, the images in these test sets cannot easily be matched to their sources or copies downloaded from other sources. To prevent the introduction of test data sourced from iNaturalist or “other sources online”, we perform two content deduplication steps. The first is to run MD5 hashes on all downloaded images. During our download, these hash sums are used to distinguish images that have already been downloaded. We record the hash of both the original image and our resized image. These were used to ensure that Rare Species [57] images (sourced from EOL) were not included in the training data.
Traditional hash sums are highly sensitive to small changes–a one-pixel difference between two images will produce a different hash. Hence, we applied perceptual hashing [PDQ 48], with distance less than , to identify training images that may be in our desired test sets that could not otherwise be filtered out (i.e., by MD5 hash sum or through metadata). Note that PDQ hash evaluation was only run on GBIF citizen science images and EOL images sourced from Flickr since they are not reliable for museum specimens.
IUCN Red List coverage. According to the most recent IUCN Red List assessment[50], TreeOfLife-200M contains images of () of all IUCN-assessed species in categories characterized as rare species or data deficient. This coverage was determined by applying our taxonomic alignment package to the IUCN taxonomic data to enable direct comparison with our standardized dataset. TreeOfLife-200M demonstrates particularly strong representation of threatened species, with coverage across threatened categories ( species), including of Vulnerable species (), of Endangered species (), and of Critically Endangered species (). Coverage extends to of Near Threatened species () and of species classified as Extinct in the Wild (). Data Deficient species have a lower representation at (), likely reflecting the challenges in imaging and identifying the species in this group. Notably, these species are designated in this category because there is not sufficient information about them for IUCN to evaluate their status; only of the M described species have been evaluated [49]. Thus, including of these species in TreeOfLife-200M, along with the threatened species coverage, establishes the approach to integrating diverse data sources used in TreeOfLife-200M as a valuable resource for conservation research, providing visual representation for a substantial majority of species prioritized for global conservation action.
Appendix J Author Contribution Statement
Jianyang Gu led the research project, ran the experiments, conducted analyses, and wrote the major part of the manuscript. For the dataset processing, he built the face detection pipeline to remove images with recognizable/identifiable people. He also retrieved the common names from GBIF backbone taxonomy alignment to TaxonoPy output.
Samuel Stevens constructed the initial benchmark for non-species classification tasks and significantly contributed to the paper writing. In addition, he also developed the content-based de-duplication pipeline based on PDQ-hash to identify test images that were contained in training data.
Elizabeth G Campolongo analyzed, evaluated, planned, organized, and documented the dataset effort. She developed the data processing approach by categories and supervised the tool development to retrieve and organize the dataset onto the HPC file system and of dataset curation. She curated camera trap test sets for benchmarking and Darwin’s Finches for embedding analysis. She also contributed significantly to the dataset part of the paper writing.
Matthew J Thompson planned, organized, and supervised the development of tools to retrieve and organize the dataset onto the HPC file system. He developed and executed the taxonomic standardization (TaxonoPy) and webdataset ML-ready format conversion. He also contributed to the paper writing of the dataset part.
Net Zhang played an essential role in cleaning the noisy data. He dealt with the museum specimen images, citizen science images, and data de-duplication. Besides, he also processed and analyzed the downloaded data, primarily focusing on GBIF.
Jiaman Wu provided detailed advice on designing the training experiments for BioCLIP 2. She also contributed to the dataset processing by initiating specimen label filtering and informing the taxonomic standardization for hemihomonym cases.
Andrei Kopanev developed the distributed-downloader tool to retrieve and organize the original images in the HPC file system and converted the dataset into ML-ready format (webdataset).
Zheda Mai provided constructive advice on designing the experience replay of CLIP training data.
Alexander E. White provided expertise in consultation on the museum specimen image processing approach.
James Balhoff and Wasila Dahdul provided expertise in consultation on the algorithm for taxonomic hierarchy resolution relevant to TaxonoPy development.
Daniel Rubenstein and Hilmar Lapp were involved in regular meetings and gave valuable feedback on results and experiments.
Tanya Berger-Wolf and Wei-Lun Chao provided insightful discussions and directions throughout the development of the project. They offered constructive advice for both the training design and the model evaluation. They also provided detailed comments on the manuscript.
Yu Su is the senior lead that oversaw the project and was involved in every aspect. He conceived the idea of observing emergent properties from scaling hierarchical contrastive training. He gave insightful guidance in formalizing the research question and directions to verify the new capabilities of BioCLIP 2. He provided critical feedback and helped shape the manuscript.

### v2

@@ Appendix G Discussion on the Relationship with Neural Collapse
First, the class prototypes in standard classification tasks and supervised contrastive learning (SCL) can both be treated as fully trainable parameters [82]. In standard classification, the prototypes are the weights of the linear classifier. In SCL, they are simply the mean embedding of the corresponding classes. The class prototypes will form a simplex ETF after extensive training. Comparatively, the adopted contrastive training scheme in BioCLIP 2 employs text embeddings of the taxonomic labels as class prototypes. During the creation of taxonomic labels, hierarchical structures have been naturally embedded into them based on ecological and functional evidence. Different species can share higher taxonomic levels, which makes them hard negatives. Therefore, even if the text encoder is also being trained, the generated prototypes will not become ETF.
@@ Appendix H Biological Visual Evaluation Details
Instead of training specialist models, we treat the evaluated models as frozen visual embedding extractors. Standard machine learning algorithms are applied on top of the acquired visual embeddings to predict the corresponding labels. Such a design is adopted to evaluate the quality of the embeddings while avoiding the influence of complicated optimization loops. In the following, we introduce the details of the adopted benchmarks and the evaluation algorithms. All the experiments are conducted with 1 NVIDIA A100 GPU, and the running time for each task is within 30 minutes.
FishNet. FishNet focuses on recognizing, locating, and predicting species and their functional traits [16]. Specifically, images are collected with annotations of habitat, ecological role, and nutritional value. In this work, we mainly focus on the prediction of habitats and ecological roles, involving 9 groups of binary labels (e.g., whether the fish can live in freshwater). Following the practice in the original paper, we train a two-layer linear classifier with binary cross-entropy loss to predict the 9 labels. We count a correct prediction only if all the 9 labels are predicted correctly for the sample. The original train-test split is adopted, where images are used in training, and the remaining images are used for testing. This task evaluates whether the embedding distribution of different species is aligned with their ecological relationships.
@@ I.1 Taxonomic Standardization
When GNVerifier returns a result from a query, our taxonomic alignment package combines it with the input taxonomic hierarchy, along with the query parameter, to form a resolution attempt. The query response, based on the most specific taxonomic term available in the input data, along with the remaining input hierarchy (entry’s resolution attempt), is then matched against pre-defined profiles. The algorithm uses these three components to fit against the series of profiles to determine whether a confident resolution is found or if an alternative query strategy is needed (such as using a different query term or data source), iterating until a match is made or alternative approaches are exhausted. If all attempts ultimately fail to match a profile, the original input data is used for the entry’s label in the final output. After hierarchical resolution, common names are annotated. For each entry, the common name corresponding to its most specific resolved taxonomic term is selected, when available, using the preferred English vernacular names from the GBIF Backbone Taxonomy [87]. In the output, the proportion of records changed for each data source is EOL: %, FathomNet: %, BIOSCAN: %, and GBIF: %. The notably low modification rate for GBIF reflects our taxonomic alignment package’s preference for the GBIF Backbone Taxonomy as its primary reference. After standardization, we summarize the image number and taxa number from each source in Figure 12. We also report the distinct taxa and image number provided by top publishers from GBIF in subsection I.1. We provide the source code for this part in TaxonoPy.
| Publisher (GBIF Data Source) |